# Generalization of Gibbs and Langevin Monte Carlo Algorithms in the Interpolation Regime

### v1

@@ 4.2 Stability of the bounds
$\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)\right]$ $\leq$ $\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[\Gamma_{\nu}\left(h,\mathbf{x},\bm{\beta}\right)\right]+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}$
$+\left(M+\beta m\right)\epsilon_{\beta}+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)m\epsilon_{\beta_{k-1}}.$
The proof, given in Section B.2, is similar to that of Theorem 3.2 and applies Markov’s inequality with $\nu_{\beta}$ instead of $G_{\beta}$. It then uses the fact that, if $f$ is a bounded measurable function, then $\left|\mathbb{E}_{\nu_{1}}\left[f\right]-\mathbb{E}_{\nu_{2}}\left[f\right]\right|\leq\left\|f\right\|_{\infty}d_{TV}\left(\nu_{1},\nu_{2}\right)$. For the single-draw version (i) this is applied to $\left(\mathbb{E}_{\nu_{\beta}}-\mathbb{E}_{G_{\beta}}\right)\left[e^{F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]$, which causes the exponential dependence on $\beta m$ and $M$. For (ii) we can apply this in the exponent to $\left(\mathbb{E}_{\nu_{\beta}}-\mathbb{E}_{G_{\beta}}\right)\big[F(h,\mathbf{x})+\beta\hat{L}(h,\mathbf{x})\big],$ and the logarithm makes the dependence linear. The rest of the proof is mechanical.
The last terms in (i) and (ii) are the additional errors due to the approximations of the Gibbs posteriors. The worst term is clearly the first one in (i) due to the exponential dependence on the proxy function $F$, which is typically of order $n$ and on $\beta$, which is larger than $n$ in the regime in which we are interested. In practice, the requirement of such an approximation is prohibitive for the single-draw version. The bound in (ii) has more moderate approximation requirements. In this case, we can also give bounds in terms of the $W_{2}$-Wasserstein metric (as guaranteed by Corollary 4.2), if $F$ sand $\ell$ satisfy a Lipschitz condition instead of boundedness. We will use the following fact: Since $W_{1}\leq W_{2}$ it follows from the Kantorovich-Rubinstein Theorem (Villani, 2009), that for any real Lipschitz function $f$ on $\mathcal{H}$ and probability measures $\nu_{1},\nu_{2}\in\mathcal{P}\left(\mathcal{H}\right)$
where $\left\|.\right\|_{\rm Lip}$ is the Lipschitz-seminorm. The following result is then immediate, with proof exactly as in (ii) of Theorem 4.3.
Theorem 4.4.
Assume the conditions of Theorem 4.3, except that instead of $\left|\ell\left(h,\mathbf{x}\right)\right|\leq m$ and $\left|F\left(h,\mathbf{x}\right)\right|\leq M$ we have $\left\|\ell\left(.,\mathbf{x}\right)\right\|_{\rm Lip}\leq m$ and $\left\|F\left(.,\mathbf{x}\right)\right\|_{\rm Lip}\leq M$ and that $W_{2}\left(\nu_{\beta_{k}}\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)=\epsilon$. Then, with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$
$\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)\right]$ $\leq$ $\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[\Gamma_{\nu}\left(h,\mathbf{x},\bm{\beta}\right)\right]+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}$
$+\left(M+\beta m\right)\epsilon_{\beta}+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)m\epsilon_{\beta_{k-1}}.$
5 Experiments
The purpose of our experiments is to show that our method gives nontrivial bounds on the test error for real-world data, while correctly bounding the test error on impossible data, where the same algorithm also achieves a small training error. For this demonstration, we use binary classification tasks. The real-world data are either the MNIST dataset, subdivided into the two classes of characters 0-4 and 5-9, or the CIFAR-10 dataset to distinguish between animals and vehicles. For impossible data, we randomize the labels of the training data. Our experiments are computationally heavy, so we generally use small sample sizes, from 2000 to 8000 examples. The hypothesis space is the set of weight vectors for a neural network with ReLU activation functions constrained by a Gaussian prior distribution with $\sigma=5$. Neural network architectures are described in Section C.1.1 of the appendix. To approximately sample the weight vectors in the vicinity of the Gibbs posterior, we use ULA as in (7) or SGLD (Welling & Teh, 2011) with constant step size $\eta$. To ensure reproducibility, we provide the code and experimental results in an anonymous repository at https://anonymous.4open.science/r/Gibbs-Generalization-45F1.
5.1 The loss function $\ell$
Unless $\sigma$ is very small, which would prevent near-zero training errors for random labels, an unbounded loss function like binary cross-entropy entropy (BCE) would cause $\mathbb{E}_{h\sim G_{0}}\big[\hat{L}(h)\big]=\mathbb{E}_{h\sim\pi}\big[\hat{L}(h)\big]$ and therefore the first term in the summation of $\Gamma$ in equation 3 to be very large. For this reason we use a bounded loss function $\ell$, either bounded binary cross-entropy as described in Appendix D of Dziugaite & Roy (2018) or the savage loss (Masnadi-Shirazi & Vasconcelos, 2008). We compute bounds for the 0-1 loss, using the method described in Section 3.3.
5.2 Approximating the ergodic mean
As we know of no sufficient criterion for convergence, we terminate iterations at time $T$, when a very slow running mean $\mathbb{M}_{\rm stop}$ of the loss trajectory $\big(\hat{L}(h_{\beta_{k}t},\mathbf{x})\big)_{t=0}^{T}$ stops decreasing. A second running mean $\mathbb{M}_{\rm erg}$ is used as an approximation of the ergodic mean and thus of expectations in the invariant distribution. We thus replace all expectations $\mathbb{E}_{h\sim G_{\beta_{k}}}\big[\hat{L}(h,\mathbf{x})\big]$ occurring in the bounds by $\mathbb{M}_{\rm erg}\big[\big(\hat{L}(h_{\beta_{k}t},\mathbf{x})\big)_{t=0}^{T}\big]$. Both running means $\mathbb{M}_{\rm stop}$ and $\mathbb{M}_{\rm erg}$ are implemented as first-order, recursive lowpass filters described in Section C.1.3 of the appendix.
5.3 Calibration
The theoretical bounds of Corollary 4.2 in combination with Theorem 4.3 can only serve as guidance for the computation of practical bounds, because the quantities $R$ and $\alpha$ are impossible to estimate in practice. But even if we assume these to be in the order of unity, the bounds are too coarse to distinguish between different temperatures with realistic stepsizes.
A simple calculation shows that $KL\left(G_{\beta},G_{2\beta}\right)\leq\beta\left(\mathbb{E}_{G_{\beta}}\left[\hat{L}\right]-\mathbb{E}_{G_{2\beta}}\left[\hat{L}\right]\right)\leq\beta\mathbb{E}_{G_{\beta}}\left[\hat{L}\right$] (Lemma B.5 in Appendix B.3). By Corollary 4.2 we should therefore have at least $8\eta dR^{2}/\alpha<\mathbb{E}_{G_{\beta}\left(\mathbf{x}\right)}\left[\hat{L}\right$] to distinguish between the expectations in the Gibbs posterior for $\beta$ and $2\beta$. The smallest neural network we use has $d=392,500$. If $\ell$ has values in $\left[0,1\right]$ then $\mathbb{E}_{G_{\beta}\left(\mathbf{x}\right)}\left[\hat{L}\right]\leq 1$, so even if $R$ and $\alpha$ are set to $1$, we would need stepsizes in the order of $10^{-7}$. Safe values of $\eta$, as suggested by the theoretical results in Section 4, are therefore impossible in practice, and the bound has to be adapted to a realistic choice of $\eta$.
The first possibility is to simply ignore the un-estimable error terms in Theorem 4.3, and to compute the bounds from $\Gamma_{\nu}\left(h,\mathbf{x},\mathbf{\beta}\right)$ in place of $\Gamma\left(h,\mathbf{x},\mathbf{\beta}\right)$ as in Theorem 3.2. We show some of these uncalibrated bounds, which already make nontrivial predictions, in Section C.2.6.
A second method uses a single calibration parameter, whose value is computed from the training data. We assume that the computed functional $\Gamma_{\nu}\left(h,\mathbf{x},\mathbf{\beta}\right)$ fails to estimate $\Gamma\left(h,\mathbf{x},\mathbf{\beta}\right)$ by a factor $r\left(\mathbf{x}\right)>0$, which we compute as
$r\left(\mathbf{x}\right)=\min\left\{r:\forall k\in\left[K\right],\text{ }\kappa^{-1}\left(\mathbb{E}_{h\sim\nu_{\beta_{k}}\left(\mathbf{x}\right)}\left[\hat{L}_{01}\left(h,\mathbf{\tilde{x}}\right)\right],\frac{1}{n}\left(r\Gamma_{\nu}\left(h,\mathbf{\tilde{x}},\mathbf{\beta}_{1}^{k}\right)+\ln\frac{2\sqrt{n}}{\delta}\right)\right)\geq\frac{1}{2}\right\},$
where $\mathbf{\tilde{x}}$ is the training set $\mathbf{x}$ with random labels and $\hat{L}_{01}$ the empirical 01-error. The calibration value $r$ is thus the smallest factor of $\Gamma_{\nu}$, for which we obtain a correct upper bound on the 01-error with random labels for all the $\beta_{k}$. By definition of $\Gamma_{\nu}$, replacing it by $r\Gamma_{\nu}$ is equivalent to a multiplicative shift of the temperature scale.
It is a purely experimental finding, that our choice of $r$ leads to correct and surprisingly tight upper bounds on the test error of correctly labeled data in all cases we tried. We emphasize that our calibration procedure depends only on the training data.
5.4 Results
Figure 1: SGLD on MNIST and CIFAR-10 with 8000 training examples, MNIST above and CIFAR-10 below, random labels on the left, correct labels on the right. Both random and true labels are trained with the same algorithm and parameters on a fully connected ReLU network with two hidden layers of 1000 and 1500 units, respectively. The calibration factor for MNIST is 0.19, for CIFAR-10 0.22. Train error, test error and our bound for the Gibbs posterior average of the 0-1 loss are plotted against $\beta$.
Several experiments confirm the validity of the proposed bounds. An example is shown in Figure 1, where a fully connected ReLU-network with two hidden layers of 1000 (respectively 1500) units each is trained with SGLD at inverse temperatures $\beta=$ 0, 500, 1000, 2000, 4000, 8000, 16000, 32000, and 64000. The train error for random labels is about 0.1 (or even less) at $\beta=64000$, where the bound is above 0.5. The test error for correct labels, however, is tightly bounded above.
Notice that for MNIST, which has the tightest bounds, the training error for the true labels is rapidly decreasing from 0.5 to 0.17 at $\beta=500$ and to 0.1 at $\beta=1000$. The more moderate initial decrease for CIFAR-10 corresponds to the tendency to overfit on this more difficult dataset. This confirms the intuition, that good generalization at low temperatures is already announced in the high temperature regime.
We generally found the uncalibrated bounds for ULA tighter than those for SGLD, consistent with the findings of Brosse et al. (2018). Experimental bounds for single draws from the posterior and various other experiments are reported in Section C.2.
6 Conclusion
Using the integral representation of the log-partition function, the Gibbs posterior admits the computation of upper bounds on the true error based on the training data and for any temperature. These bounds are stable under perturbation in the total-variation and Wasserstein metrics, and can be approximated by Langevin Monte Carlo (LMC) algorithms. However, for realistic experiments, the approximations obtained by these algorithms are coarse and require calibration, which leads to rather tight bounds in the interpolation regime of overparametrized neural networks.
Acknowledgments
This work was supported in part from the PNRR MUR Project PE000013 CUP J53C22003010006 "Future Artificial Intelligence Research (FAIR)", funded by the European Union – NextGenerationEU and by European Lighthouse of AI for Sustainability, Grant N. 101120237.
Appendix
In this appendix, we summarize a glossary of notation, give additional theoretical results and missing proofs, and provide more information on the numerical experiments, as well as additional experimental results.
Appendix A Table of notation
Notation Brief description Section
$\mathcal{X}$ space of data 2
$\Sigma$ sigma algebra (events) on $\mathcal{X}$ 2
$\mu$ probability of data 2
$n$ sample size 1, 2
$\mathbf{x}$ generic member $\left(x_{1},...,x_{n}\right)\in\mathcal{X}^{n}$, training sample 2
$\mathcal{H}$ hypothesis space 2
$\Omega$ sigma algebra (events) on $\mathcal{H}$ 2
$\ell$ $\ell:\mathcal{H\times X\rightarrow}\left[0,\infty\right)$ loss function 2
$f$ secondary loss function 3.3
$\mathcal{P}\left(\mathcal{H}\right)$ probability measures on $\mathcal{H}$ 2
$\pi$ nonnegative a-priori measure on $\mathcal{H}$ 2
$\sigma$ width of Gaussian prior 4
$L\left(h\right)$ $L\left(h\right)=\mathbb{E}_{x\sim\mu}\left[h\left(x\right)\right]$, expected loss of $h\in\mathcal{H}$ 2
$\hat{L}\left(h,\mathbf{x}\right)$ $\hat{L}\left(h,\mathbf{x}\right)=\left(1/n\right)\sum_{i=1}^{n}\ell\left(h,x_{i}\right)$, empirical loss of $h\in\mathcal{H}$ 2
$\beta$ inverse temperature 1, 2
$Z_{\beta}\left(\mathbf{x}\right)$ partition function 2
$G_{\beta,\pi}\left(\mathbf{x}\right)$ Gibbs posterior with energy $\hat{L}$ and prior $\pi$ 2
$\mathbb{E}_{g\sim G_{\beta}\left(\mathbf{x}\right)}$ posterior expectation 2
$\bm{\beta}$ increasing sequence $\left(\beta_{1}<...<\beta_{K}\right)$ of positive reals 3.1
$\Gamma\left(h,\mathbf{x},\bm{\beta}\right)$ bounding functional 3.1
$F\left(h,\mathbf{x}\right)$ placeholder for generalization gap 3.2
$\kappa$ $kl\left(p,q\right)=p\ln\frac{p}{q}+\left(1-p\right)\ln\frac{1-p}{1-q}$, rel. entropy of Bernoulli variables 2
$KL\left(\rho,\nu\right)$ $\int\left(\ln\frac{d\rho}{d\nu}\right)d\rho$, KL-divergence of $\rho,\nu\in\mathcal{P}\left(\mathcal{H}\right)$ 2, 4.1, 4.2
$d_{TV}\left(\rho,\nu\right)$ total variation distance 2, 4.1, 4.2
$W_{p}\left(\rho,\nu\right)$ $p$-Wasserstein distance 2, 4.1, 4.2
$\Gamma_{\nu}\left(h,\mathbf{x},\bm{\beta}\right)$ LMC approximation of $\Gamma\left(h,\mathbf{x},\bm{\beta}\right)$ 4.2
$\eta$ step size or learning rate 4.1
$\nu_{\beta,\eta}$ invariant measure of LMC approximation of $G_{\beta}$ with step size $\eta$ 4.1
$\nu_{\beta,\eta,t}$ LMC approximation of $G_{\beta}$ with step size $\eta$ at iteration $t$ 4.1
$r({\mathbf{x}})$ calibration factor 5.3
$\mathbf{\tilde{x}}$ randomly labeled data 5.3
$\mathbb{M}_{\rm stop}$$\mathbb{M}_{\rm erg}$ filters for stopping and ergodic mean 5.2, C.1.3
Appendix B Additional results and proofs
B.1 Sub-gaussian losses
The freedom in the choice of $F$ allows a number of bounds to be derived from Theorem 3.2. A centered real random variable $Y$ is called $\sigma$-sub-Gaussian if $\ln\mathbb{E}e^{\lambda Y-\mathbb{E}Y}\leq\lambda^{2}\sigma^{2}/2$ for all $\lambda\in\mathbb{R}$. Now, suppose that for some real function $f$ all the $x\in\mathcal{X\mapsto}f\left(h,x\right)$ are $\sigma$-sub-Gaussian as $x\sim\mu$. Let $\hat{L}(h,\mathbf{x})=\frac{1}{n}\sum_{i=1}^{n}f(h,x_{i})$ and $L(h)=E_{x}[f(h,\mathbf{x})]$. Then, $\mathbf{x}\in\mathcal{X}^{n}\mapsto\hat{L}\left(h,\mathbf{x}\right)$ as $\mathbf{x}\sim\mu^{n}$ is $\sigma/\sqrt{n}$-sub-Gaussian. It is tempting to set $F\left(h,\mathbf{x}\right)=\lambda\left(L\left(h\right)-\hat{L}\left(h,\mathbf{x}\right)\right)$ in Theorem 3.2, divide by $\lambda$ and then optimize over $\lambda$. Unfortunately, the last step is impossible, since the optimal $\lambda$ is data-dependent in its dependence on $\Gamma$ and ruins the exponential moment bound on $F$. A more careful argument establishes the following.
Corollary B.1.
Suppose that for all $h\in\mathcal{H}$ the random variables $x\in\mathcal{X\mapsto}f\left(h,x\right)$ as $x\sim\mu$ are $\sigma$-sub-Gaussian. For $\delta>0$ with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$ and $h\sim G_{\beta,\pi}\left(\mathbf{x}\right)$ if $\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\geq 1$, then,
$E_{x}\left[f\left(h,x\right)\right]-\frac{1}{n}\sum_{i}f\left(h,x_{i}\right)\leq\sigma\sqrt{\frac{2[\Gamma\left(h,\mathbf{x},\mathbf{\beta}\right)\left(1+1/n\right)+\ln\left(\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\left(n+1\right)/\delta\right)]}{n}}.$
For the proof, we use the following auxiliary result.
Lemma B.2.
(Anthony & Bartlett, 1999, Lemma 15.6) Suppose $\Pr$ is a probability distribution and
$\left\{E\left(\alpha_{1},\alpha_{2},\delta\right):0<\alpha_{1},\alpha_{2},\delta\leq 1\right\}$
is a set of events, such that
(i) For all $0<\alpha\leq 1$ and $0<\delta\leq 1$,
$\Pr\left\{E\left(\alpha,\alpha,\delta\right)\right\}\leq\delta.$
(ii) For all $0<\alpha_{1}\leq\alpha\leq\alpha_{2}\leq 1$ and $0<\delta_{1}\leq\delta\leq 1$
$E\left(\alpha_{1},\alpha_{2},\delta_{1}\right)\subseteq E\left(\alpha,\alpha,\delta\right).$
Then for $0<a,\delta<1$,
$\Pr\bigcup_{\alpha\in\left(0,1\right]}E\left(\alpha a,\alpha,\delta\alpha\left(1-a\right)\right)\leq\delta.$
We put this lemma in a more convenient form.
Lemma B.3.
Let $Y$ and $X\geq 0$ be real random variables, $\psi:\mathbb{R}\times\left(0,1\right)\rightarrow\mathbb{R}$ be increasing in the 1st argument and $\forall C>1$, $\delta\in\left(0,1\right)$,
$\Pr\left\{X\leq C\wedge Y>\psi\left(C,\delta\right)\right\}<\delta.$
Then, for every $\epsilon>0$
$\Pr\left\{X\geq 1\wedge Y>\psi\left(X\left(1+\epsilon\right),\frac{\delta\epsilon}{X\left(1+\epsilon\right)}\right)\right\}.$
@@ B.1 Sub-gaussian losses
This follows from Lemma B.2 using the events
$E\left(\alpha_{1},\alpha_{2},\delta\right)=\left\{X\leq\alpha_{2}^{-1}\wedge Y>f\left(\alpha_{1}^{-1},\delta\right)\right\}$
and $a=1/\left(1+\epsilon\right)$. ∎
Proof of Corollary B.1.
Take $F=\lambda\left(L\left(h\right)-\hat{L}\left(h,\mathbf{x}\right)\right)$, so $\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{g\sim\pi}\left[e^{F\left(g,\mathbf{x}\right)}\right]\leq\lambda^{2}\sigma^{2}/\left(2n\right)$. From Theorem 3.2 and the properties of sub-Gaussian variables we get with $\lambda=\sigma^{-1}\sqrt{2n\left(C+\ln\left(1/\delta\right)\right)}$ that
$\Pr\left\{\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\leq C\wedge L\left(h\right)-\hat{L}\left(h,\mathbf{x}\right)>\sigma\sqrt{\frac{2\left(C+\ln\left(1/\delta\right)\right)}{n}}\right\}$
$=$ $\Pr\left\{\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\leq C\wedge L\left(h\right)-\hat{L}\left(h,\mathbf{x}\right)>\frac{C+\ln\left(1/\delta\right)}{\lambda}+\frac{\lambda\sigma^{2}}{2n}\right\}$
$=$ $\Pr\left\{\Gamma\left(h,\mathbf{x},\bm{\beta}\right)\leq C\wedge\lambda\left(L\left(h\right)-\hat{L}\left(h,\mathbf{x}\right)\right)>C+\lambda^{2}\sigma^{2}/\left(2n\right)+\ln\left(1/\delta\right)\right\}\leq\delta.$
Substitution in Lemma B.3 with $\psi\left(C,\delta\right)=\sigma\sqrt{2\left(C+\ln\left(1/\delta\right)\right)/n}$, $X=\Gamma\left(h,\mathbf{x},\bm{\beta}\right)$, $Y=L\left(h\right)-\hat{L}\left(h,\mathbf{x}\right)$ and $\epsilon=1/n$ gives Corollary B.1. ∎
B.2 Proofs for Section 4.2
Restatement of Theorem 4.3:
Theorem B.4.
Suppose that $\mathcal{H}=\mathbb{R}^{d}$ and that there are numbers $m,M<\infty$ such that for every $\mathbf{x}$ in $\mathcal{X}^{n}$ and $h\in\mathcal{H}$ we have $\left|\ell\left(h,\mathbf{x}\right)\right|\leq m$ and $\left|F\left(h,\mathbf{x}\right)\right|\leq M$. Let $0=\beta_{0}<\beta_{1}<\cdots<\beta_{K}=\beta$ and $\nu_{\beta_{k}}\left(\mathbf{x}\right)\in\mathcal{P}\left(\mathcal{H}\right)$ be such that $d_{TV}\left(\nu_{\beta_{k}}\left(\mathbf{x}\right),G_{\beta_{k}}\left(\mathbf{x}\right)\right)=\epsilon_{\beta_{k}}$. Then,
(i) with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$ and $h\sim\nu_{\beta}\left(\mathbf{x}\right)$
$F\left(h,\mathbf{x}\right)$ $\leq$ $\Gamma_{\nu}\left(h,\mathbf{x},\mathbf{\beta}\right)+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}$
$+\ln\left(2e^{M+\beta m}\epsilon_{\beta}\right)+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)m\epsilon_{\beta_{k-1}}.$
(ii) with probability at least $1-\delta$ as $\mathbf{x}\sim\mu^{n}$
$\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[F\left(h,\mathbf{x}\right)\right]$ $\leq$ $\mathbb{E}_{h\sim\nu_{\beta}\left(\mathbf{x}\right)}\left[\Gamma_{\nu}\left(h,\mathbf{x},\bm{\beta}\right)\right]+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}$
$+\left(M+\beta m\right)\epsilon_{\beta}+\sum_{k=1}^{K}\left(\beta_{k}-\beta_{k-1}\right)m\epsilon_{\beta_{k-1}}.$

### v2

@@ C.3 The "Second law of thermodynamics"
$\mathbb{E}_{\nu\left(\mathbf{x}\right)}\!\left[F(h,\mathbf{x})\right]$ $\hskip-15.6491pt\leq\hskip-14.22636pt$ $\hskip-2.84544pt-\beta\mathbb{E}_{\nu\left(\mathbf{x}\right)}\big[\hat{L}\left(h,\mathbf{x}\right)\big]+\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})+\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right]+\ln\frac{1}{\delta}$
$+KL\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)+\Delta.$
@@
Proof of (ii). By equation (6) and Lemma 4.4 we have
$KL\left(\nu\left(\mathbf{x}\right),\pi\right)$ $\leq$ $KL\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)-\beta\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]+\int_{0}^{\beta}\mathbb{E}_{h\sim G_{\gamma}\left(\mathbf{x}\right)}d\gamma$
$\leq$ $KL\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)-\beta\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[\hat{L}\left(h,\mathbf{x}\right)\right]+\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})+\Delta.$
Substitution in (ii) of the PAC-Bayesian theorem proves (ii).
Proof of (i) without amendment. Similarly, we have from Lemma 4.4
$\ln\frac{d\nu\left(\mathbf{x}\right)}{d\pi}\left(h\right)$ $=$ $\ln\frac{d\nu\left(\mathbf{x}\right)}{dG_{\beta}\left(\mathbf{x}\right)}\left(h\right)+\ln\frac{dG_{\beta}\left(\mathbf{x}\right)}{d\pi}\left(h\right)$
$\leq$ $R_{\infty}\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)-\beta\hat{L}\left(h,\mathbf{x}\right)+\int_{0}^{\beta}\mathbb{E}_{h\sim G_{\gamma}\left(\mathbf{x}\right)}d\gamma$
$=$ $R_{\infty}\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)-\beta\hat{L}\left(h,\mathbf{x}\right)+\Gamma(\nu_{0}^{K-1},\mathbf{x},\beta_{0}^{K})+\Delta.$
Then substitute in the PAC-Bayesian theorem (i).
For the proof of the amendment to (i) we have to return to the proof of the PAC-Bayesian theorem. From Markov’s inequality we have (in analogy to the proof of Theorem 2.1) with probability at least $1-\delta$ as $x\sim\mu^{n}$ and $h\sim\nu\left(\mathbf{x}\right)$, that
$\leq$ $\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\nu\left(\mathbf{x}\right)}\left[e^{F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]+\ln\left(1/\delta\right)$
$\leq$ $\ln\left(\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim G_{\beta}\left(\mathbf{x}\right)}\left[e^{F\left(h,\mathbf{x}\right)+\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)}\right]+e^{\beta\left\|\ell\right\|_{\infty}+\left\|F\right\|_{\infty}}d_{TV}\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)\right)+\ln\left(1/\delta\right)$
$=$ $\ln\left(\max\left\{\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right],1\right\}+\max\left\{e^{\beta\left\|\ell\right\|_{\infty}+\left\|F\right\|_{\infty}}d_{TV}\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right),1\right\}\right)+\ln\left(1/\delta\right)$
$\leq$ $\max\left\{\ln\mathbb{E}_{\mathbf{x}}\mathbb{E}_{h\sim\pi}\left[e^{F\left(h,\mathbf{x}\right)}\right],0\right\}+\max\left\{\beta\left\|\ell\right\|_{\infty}+\left\|F\right\|_{\infty}+\ln\left(2d_{TV}\left(\nu\left(\mathbf{x}\right),G_{\beta}\left(\mathbf{x}\right)\right)\right),0\right\}$
$\ \ \ +\ln\left(1/\delta\right).$
In the second inequality we used $\ln Z_{\beta}\left(\mathbf{x}\right)\leq 0$ and the property of the total variation metric. In the next inequality we used Lemma 3.1, and in the last line we used for $a,b\geq 1$ that $\ln\left(a+b\right)\leq\ln\max\left\{a,b\right\}+\ln 2\leq\ln a+\ln b+\ln 2$. Subtract $\beta\hat{L}\left(h,\mathbf{x}\right)+\ln Z_{\beta}\left(\mathbf{x}\right)$, use Lemma 3.1, Lemma 4.4 and Pinsker’s inequality to bound the total variation distance.
C.5 Miscellaneous Lemmata
Lemma C.7.
For $0<\beta<\infty$
$\max\left\{KL\left(G_{\beta},G_{2\beta}\right),KL\left(G_{2\beta},G_{2\beta}\right)\right\}\leq\beta\left(\mathbb{E}_{h\sim G_{\beta}}\left[\hat{L}\left(h\right)\right]-\mathbb{E}_{h\sim G_{2\beta}}\left[\hat{L}\left(h\right)\right]\right)$